OpenStudy (anonymous):

What is the y coordinate of the y - intercept of the line that passes through the points (-2,-2) and (2,4) ?

OpenStudy (anonymous):

@campbell_st could u help me?

OpenStudy (campbell_st):

ok... the slope of the line is m = (4 - -2)/(2 - - 2) m = 6/4 or 3/2 so the lin is in the form y = 3/2 x + b substitute either point into the above equation to find b... the y-intercept

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

I so it's one !! Thanks @campbell_st